OpenStudy (anonymous):

the westside factory produces 450 boxes of cereal everyday. a macjine drops a single toy car into each box. today the machine is loaded with 115 blue cars, 190 green cars, and 145 red cars. what is the probability that one of the boxes of cereal has a blue car inside

OpenStudy (gw2011):

Since there is a total of 450 boxes and cars and 115 blue cars, then the probability should be 115/450 = 23/90 = 0.2555 which is approximately a little over 1/4--that is, a little over 1 out 4 probability.